OpenStudy (anonymous):

kim609 Up 1 Down Topic: Math Find the value of x that satisfies each equation. A) 5 1/6x+x=8

OpenStudy (anonymous):

x-1/4 = 3/4` To solve we must get x by itself so add `1/4` to both sides of the equation: `x-1/4+1/4 = 3/4 +1/4` `therefore x= 4/4`